FEL Upgrade August 2002 Monthly Report

There has been much progress in installation around the machine and fabrication of magnets and vacuum components (although much remains). The 3F region of beamline, quadrupole girders (downstream of the third cryomodule before the bend) was connected by the vacuum group.

Three more GWs were received from the vendor which completes the group required for the IR machine. They will be characterized in the magnet test stand before installation. Virtually all of the remaining stands were installed in the tunnel and substantial progress is now being made on installation now that full crews have returned to the FEL from the main machine maintenance period.

One of the test/assembly stands was set up with a complete deformable mirror assembly, actuators and LVDT. This will permit testing of the alignment accuracy and response frequency with the full mass driven.

Work on the crucial ball cathode support tube was completed this week. See the link below for an image of the prepared components. The high voltage power supply for the gun also made significant assembly progress.

A technical requirements and preliminary design review was held on Aug. 22 on the laser microengineering experimental station that is being designed and built for the FEL Upgrade by the Aerospace Corporation as part of our joint contract with the AFRL. The review (which was held by teleconference to save travel time and costs) was judged to be successful.

We are commencing the Accelerator Readiness Review Process in anticipation of turning on the laser. See the web link below.

On August 8 we briefed Mr. Peter Teets, the Under SECAF(Space/Director, NRO) and his Deputy for Military Space, Robert Dickman on the Navy scale-up program.

We thank the Office of Naval Research for sending us an additional $80k of the $300k remaining FY02 allocation for the agreed tasks.

H. F. Dylla spent 2.5 days at a Director’s Council offsite working out-year program planning and priorities for the laboratory.

We thank the ONR for their continuing efforts to keep our accounts in the black by completing delivery (plus some) of the remaining FY02 funds due. On Aug. 19-20, Jefferson Lab held a review of the lab's Institutional plan for the new head of DOE's Office of Science, Ray Orbach. The OMS/AFRL funded FEL project was included in both the formal presentation and site tour.

The project technical and financial report for the month of July was completed and distributed to the DOE, ONR and AFRL contract monitors.

On Aug. 23, F. Dylla and G. Neil attended a briefing to the Deputy Secretary of the Navy (Livingstone) presented on behalf of Admiral Cohen (ONR) and Mathis (NAVSEA).

The dates for the next semiannual review of the FEL Upgrade project has been selected by the review committee: Nov. 20-21.

George Neil attended a Pulsed Laser Lethality Workshop and presented a description of our Upgrade and schedule to allow interested parties to plan on utilizing the laser for materials tests. Discussions were also held with a group from LANL concerning the possibility of testing predictions that at sufficient peak intensities compensation of thermal blooming can be accomplished by using non-linear effects of the atmosphere. It appears the IR Upgrade is in the proper parameter range to test these ideas.

Following the Workshop a separate meeting of the JTO HEL Lethality TAWG was held. Discussions on prioritization of tasks for the coming years included establishing a database of prior measurements and (on the Navy side) planning for establishment of test facility at JLab. A report detailing this roadmap is in preparation for November to present to the JTO.

The internal JLab web-based FEL ARR (Accelerator Readiness Review) process is up and running. Sub-systems managers are using it to document their sub-system's Readiness status and to build an ARR reference documents database. The following is a link to the main ARR Systems Status Home Page. (Note: For security purposes, the link and database is behind the Jlab firewall.) http://devsrv-e.acc.jlab.org/safety/fel/index.php

Project Cost Performance:
The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $62k were accrued for August for total accrued costs of $4,080k to date for Phase 2. Work scheduled for August was $418k for a total of $3,937k for Phase 2 to date; work performed for August was $289k for a total of $3,803k for Phase 2 to date, which results in 89% scheduled vs. 86% performed for the Phase 2 project. At present the cost variance for Phase 2 is -$276k and the schedule variance is -$134k. We expect to meet the planned cost at completion of Phase 2 effort. The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project (Phase 1) was approved for $2,836k for a performance period from September 2001 to September 30, 2002. Actual charges of $126k were accrued through August for a total accrued costs of $2,069k since project start. Work scheduled for August was $95k for a total of $2,231k for the UV project to date, work performed for August was $502k for a total of $2,395k for the UV project to date, which results in 79% scheduled vs. 84% performed for the UV project. At present the cost variance for the UV project is $325k and the schedule variance is $164k. The program is on track to meet the statement of work requirements on schedule and within the available budget.

Phase 2 of the UV FEL Project was approved for $1,472k for a performance period from August 2, 2002 to September 30, 2003. Expenses to date are $381k with commitments of $147k.

WBS 3 (Beam Physics):
A pi-mode based n-cell cavity model for user-defined RF wavelength was constructed. Beamlette stacking was investigated.

In response to an inquiry by S. Benson, and with information he kindly provided, a driver configuration supporting the 1 micron FEL option was developed. It is identical to the rev113 design save for an inward move (toward the wiggler) of the quad triplets adjacent to the FEL insertion. This was required to avoid generating large beam envelopes (with associated aberrations) where the small, strongly divergent beam enters/exits the shorter wiggler.

WBS 4 (Injector):
The SF6 aluminum tank is ready for shipment to be modified. The fixtures for welding the support tube to its mounting flange are being fabricated in the machine shop. The support tube was e-beam welded to its flange. The rear stalk bellows drive unit design is in review. The gun oven extension drawings are in review. All off the internal hardware for assembly of the gun has been vented and is out to be silver plated. Drive hardware for the running resistor is on order. The fixturing for welding the support tube to its flange was fabricated.

The work with W&M on the ion implantation of the ball cathode support tube concluded successfully this week. A close examination revealed that the tube is indeed free of arcing marks. A total of six runs were performed to cover the entire surface of the tube with SiO2.

A SS sample will be implanted and field emission scanned for testing up to 60 MV/m (the highest field region on the tube will be about 12 MV/m). Pictures of the implanted support tube and the ball cathode can be seen at http://laser.jlab.org/gun/.

POISSON simulations of the two gun solenoids have been performed to study the effect of the increased aperture on the downstream solenoid. It is important to verify this because PARMELA input files have not been updated with such modification. So far, the simulations show a variation no larger than 2% in the peak magnetic field. A more detailed analysis is in progress.

Gun HVPS - The 3 stacks and their toroids are assembled in the HVPS Tank. The electronics is ready to mount in their holders on the side of the tank. The piping for the SF6 to the tank is nearly complete. Drawings for the running resistor housing were signed off and a new Conditioning Resistor was made. Parts for the combined resistors and transmission line are being made by the Machine Shop.

WBS 5 (SRF):
Studies of HOM damping in the 7 cell cavities continue including Q measurements on the actual cavities from the first (CEBAF) module and calculations of stability limits using TDBBU and other codes. All 8 cavities of this first CEBAF module are to be passed over for string assembly next week.

Plans to measure HOM damping in a 7 cell cavity continue. In order to get realistic measurements of the HOM Qs we intend to test a seven cell cavity at temperature in a cryogenic dewar with both the high power and HOM couplers attached. This test can be initiated within two weeks or less after the required dewar fit up hardware is machined. Copper cavity tests are also continuing to determine required placement of the HOM loads for optimal coupling.

Preparations continue and are nearly complete for measurements of a full cavity HOM damping Qs including high power coupler effects. Modifications to the vertical dewar mounting were readied. In addition estimates of beam breakup thresholds were calculated based on earlier measured values as well as copper model measurements.

WBS 6 (RF):
Quarter HVPS - Continued debugging the hardware and P/C boards. Initial talks were held with G. Lahti to get started on the communication link between the PLC in the HVPS and EPICS for remote operation and display. The waveguide was connected for unit 3 to enable No-Load testing of the HVPS. The interlocks have all been checked. After Unit 3 was hi-potted to 40 kV, the 480 VAC was connected. The unit operated to 27 kV under no-load conditions which is the maximum voltage available from the 70% tap position. We have experienced faulty klystron body current, arc, and crowbar trips that occur when the heater relay makes an adjustment. It appears to be caused by the +24 V power supply being overloaded. A power supply with more current capacity is being tried. Work continued on the Crowbar and grounding system for the HVPS. The thyratron for the Unit 4 was tested and found to be faulty. The 2 old faulty thyratrons are being returned to the vendor for estimates to restore them to working order. A P.O. has been placed to replace the thyratron we borrowed and tested from the SNS project. Progress is being made on the grounding system, but 1.5k amps with nanosecond rise time does make healthy transients.

Quarter Klystrons -- The filament resistance of the 100 kW klystrons appear to be slowly increasing. A. Mizuhara is not concerned at this time. We will continue to monitor the klystron now that the instrumentation is settling down. Hoses were fabricated this week for the Circulator, RF Load, and HOM LCW manifolds.

Zone 4 - The RF waveguide connections to the cryomodule were made by SRF this month. The balance of the waveguides will be connected when the parts are available. Flexible couplings are on order and the half-height sections are being made by the Machine Shop.

Zone 2 - The waveguide sweeps are ready for installation. The cable pulls past this zone is complete.

Zone 1 - J. Preble is scheduling the waveguide connections to the Quarter in preparation for SRF commissioning. The warm windows have been changed for higher RF power.

WBS 8 (Instrumentation):
Drive Laser Pulse Control (DLPC) - A 64bit digital I/O card (GPVS64) has been installed into iocfel10. This board will provide the necessary control for the CO305 card (for micropulse control) and the basic user requests to open/insert and close/remove the MPS shutter and the ND2 filter respectively. The new EPICS screen for rep-rate and pulse width control have gone through the first wave of testing. It is still in the development stage and is currently being revised. Screen-shot at: http://laser/systems/dlpc/upgrade2002/dlpc-newscreen-rev1.jpg. Progress on the PCB layout for the CO-305 synchronous countdown (micropulse selection) carrier board continues as well. The new implementation of this device will provide the full capability of the con-optics into EPICS.

Beam Loss Monitors (BLMs) - The fabrication drawings for the BLM front board (F0151) have been signed-off and the board was sent out for prototype production. The majority of the parts are in-house and ready for assembly. The assembly and testing will be done in stages. The BLM rear board (F0150) is being ECO'ed to correct a component placement issue in regards to mating with the cable from the BLM heads. Additionally, the requirements for the software have been worked out as well and will be ready as the F0151board goes into its testing phase. Issues surrounding the software upgrade to the high voltage power supply (LeCroy) and related screens have been addressed the new screens are under development. 1 of 6 of the BLM front boards (F0151) has been partially populated and is undergoing some simple performance tests. Upon completion of the testing the board will be submitted back to have the next group of components installed. Everything is proceeding well. The work for ECO of the BLM rear cards (F0150) has been completed, however, the updated drawing packages have not been delivered as yet.

The DSP/Pyroline camera project continues to make progress. This week we were able to output the dummy_clk_in from a timer (pretty much a noise free signal) and also able to get the dummy_pulse out of a GPIO port. Combined the two routines (a) generation of dummy_clk_in signal and (b) outputting a dummy_pulse waveform upon feeding the DSP with dummy_beam_sync. Testing has been good so far. Upon request from user (that's just temporary), A/D conversion can be achieved. We've been working on the part of inserting A/D with the rest of the 2 routines. If this starts to work then the timing issues need to be tended to next.

Personel Safety System (PSS)-Completed the cable installation for the Magnet Personel Safety Interlock. Terminations and checkout are still in progress. The PSS RF interface cables have been terminated into the PSS PLC's and into their interface chassis in Zone 4. Work has been completed on the assembly of the super trim racks in Zones 1 and 5. Klixon cables have been pulled into their new destinations. The BPM cables have dressed in and labeled in Zone 4.

The DSP code is progressing for the spectrometer (and to be used for all other apps. Like BPM, power supplies...). This month’s effort by our student was on running the routine for outputting a waveform (serves to form the pulse) from a GPIO pin when a signal is fed into one of the timer pins (serves to form the beam sync). The signal received from the GPIO is a bit noisy (either a glitch in the code or needs filtering, hence executing code that would remove the noisy component from the signal(this will prove helpful in the future as well).

MPS - Trent Allison from EEI&C is helping out on the design of the Machine Mode/Beam Mode generator cards. In discussions this week concerning the layout and packaging of the system were are able to proceed with a design that does not need two separate boards for these functions. A 25-pin "D" ribbon cable will straddle the input boards and the gate arrays for each board will be mapped to all of the pins (20 not the 5 ground pins) such that each board will have three pins to 'talk' on and will 'listen' to the rest. The boards will be in-circuit programmable via a 10 pin header on the front panel. The documentation for the 6U (VME) isolation board (GPIM64/05516-B-0009) has been collected and the components needed to assemble 4 to 6 of these cards are being procured.

The design review was held for the large magnet power supply installation and switch configuration there were very minor comments - good job done by the EEDC group, thanks! The old CAMAC crate and electronics for the magnet interface have been removed from Zone 5 to facilitate the effort on the magnet communication interface upgrade. This leaves only the RF system in CAMAC. The RS-485 communication cables have been pulled into place for the trim magnet utility chassis. System and installation drawings are being updated as work progresses for this system. Cables were pulled in Zone 5 for the Magnet Thermal Interlock chassis. The 4000 feet of 2 ga. cable is here for inject, extract, and octa-sextapole magnets, the student will enjoy pulling it in over the weekend.

Further progress has been made with the PC/104 board. The students were able to get EPICS to compile correctly under linux on the desktop, and compile an example application (i.e. was able to get an epics IOC shell prompt on the PC). The example code (or any application that we develop) can be copied directly off the PC to the PC/104 board and will run without modification. The PC/104 will serve as a low (~no) cost alternative to the VME based data acquisition solution. The first application is to monitor the optical cavity mirror thermocouples.

Great progress has also been made with the automated emittance measurement project. A new video capture card was selected that will allow maximum usage of the Dell PC for video processing. The one we were using is already obsolete - shelf life worse that fruit. The new card is an Osprey-100. It has three composite video inputs and one s-video input on a single PCI card. The spot size analysis code is being reworked to use multidimensional array, rather than a single dimension array. This will aid in data manipulation and analysis. The image captured is now moved into a multidimensional array, and can be saved. The summation and deletion (background subtraction) functions will be patched up to work the same for consistency. Some time was spent figuring out how to properly pass a multidimensional array with C. This design could also be used for control of the optical cavity mirror heaters. These mirror heaters are to be controlled base on a HeNe spot size.

A trip was made to Fermi Lab this month to review their new switched mode trim power supplies. They have developed a range of four quadrant power converter modules that cover the range of + 10 Amp, + 30 Amp, + 50 Amp, and + 100 Amp (2 50's in parallel). These are 7" high modules that can go 8 (10 Amp) across, the approach uses a single bulk supply at 150 Volts and FET switchers at 40KHz. We are getting a quote from a commercial design shop for adding a DSP/CAN interface to the control card.

The Fermi Lab trim card design was taken to a design engineering company for a quote on modifying the control card to hold a DSP. This has the potential to satisfy all of the FEL's DC Power needs from the 1 Amp trims to the 100 Amp inject and extract dipole strings. Work has begun on the 8 channels of floating power supplies for the arc quad windings. These will boost/buck fields in a larger magnet to produce the required quad fields. Installation of the DB/DJ correctors is in progress for the 2F and 3F region. Completed cable removal for the old magnet shunts. This is the last large system targeted for cleanup/ removal to make room for new installation.

Design and fabrication of the gun High Voltage power supply interface is progressing. The new VME interface (Acromag card)channel assignments have been made are being programmed by Al. The PCB prototype for this system is nearing completion. The optical cavity wiring design is complete, Al is writing the software now. The interface chasssis design is complete but there is no manpower to start it yet.

The RS-485 cables have been terminated for the trim racks. Documentation is in the final stages for the 64 bit VME isolation daughter module for the MPS system. Power cables (535 MCM) were pulled from the 66 kW box supply to the to the wiggler, and 2 ga. cable has been pulled to sex/octapole magnet locations. Beam line has been installed in the dummy cryomodule section of the linac.

Further work has been completed on the Linux spot size code. The "background deletion" function now works with a two dimensional array, and the image capture and save functions have been debugged. Our current estimate on the performance of this program is 5-7 frames-per-second (fps), though we believe that can be improved through optimization of the capture code(the analysis functions are reasonably efficient). We will be getting in contact with Al Grippo soon to start implementing the EPICS part of this project. Code will be written to detect the tick marks, beam spot, and estimate the spot size, distribution, and position.

MISC - some much needed maintenance was done on our data/web server (i.e. laser.jlab.org) The insertable dump water interlock requirements were identified and are being implemented by the mechanical support group.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
•  All coils are potted.
•  We received the next three GWs.
•  The first GW underwent rudimentary magnet field tests for core flatness on the auxiliary magnet test stand. First data indicates
    core field is near the plus or minus 1 part in 10000 flatness over the good field region.
•   The second GW underwent magnet field tests for core flatness on the primary magnet test stand. As with the first magnet, data
    indicates core field is near the plus or minus 1 part in 10000 flatness over the good field region.
•   The remaining 6 cores of the second batch of 6 Magnets for UV had ambiguous gap dimension readings at Diamond Tool, the
    subcontract machining house. Though the maximum spread in the readings is at the .003 inch level, (a hair thickness) our spec is
    one fifth of that. MEI is orchestrating the measurements and decisions on how the parts are to be brought into specification
    through selective grinding
Injector Dipoles (DU/DV)
•   The GVs were assembled and the first and second were shipped and received by us.
•   We resolved some contract issues with Wang NMR for work outside of the scope contract.
•   The first GV started magnet measurements on the magnet test stand The measurements of the first GV verified that the scheme of biasing the core field with a trim coil allows the central field to be made flat to within ± one part in 10,000, a very difficult feat for this "all ends" magnet.
• The field flatness comes with a price however. The main current is 9 A higher than the field in the GU partner magnet in the string. The GU will require a shunt that had not been anticipated.
• In addition, both the field flatness and central field are very dependent on trim coil ramp rate.
• The first try at finding the main and trim current at various field levels corresponding to various injection energies was tried. The currents are almost linear and can be extrapolated to intermediate energies.
• The first integrals were taken.
• Our next concentration is the value of the integrals on either side of the centerline of this wedge magnet.
• Repeatability is still a problem. We are working on hysteresis and power supply qualities.
• The gluing of shim materials to GU cores passed after examining the detailed records o of the individual measurements and GU magnets were assembled.
• The first GU arrived by air freight and look good.
Arc 180 Degree Dipoles (GY)
• At Wang NMR, the winding of the fourth GY Coil is complete.
• GY Coil Potting fixture is almost ready.
• We worked out the contractual issues associated with winding a fifth coil to replace the potentially unreliable first coil.
• The conductor for the fifth coil is in insulation preparation.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• Process Equipment Co. (PECo) assembled the remaining three GQ cores and shipped them and Wang NMR received them.
• At Wang NMR, potting fixtures remain in fabrication for the GQ and GX coils.
• All GX and GQ coils are wound.
• The GX cores have undergone gap inspection and are of similar high quality to the GUs.
Multipole Magnets
Trim Quad (QT)
• The first unit passed temperature test with added cooling of the side coils.
• The first unit was tested using the movable traverse method on the quad test stand. The data is being analyzed.
• Milhous should be shipping four trim quads with the auxiliary cooling plates on the side coils. We are ready to test.
Sextupole (SF)
• We have decided to not use the 6.6 inch thick plate we originally bought for material for the Sextupole project. Rather, we will use the much thinner spare plates that were bought for the dipole cores to obtain the 4 inch thick pieces. This plan eliminates a lot of "hogging" machining. It also uses up a batch of disparate pieces (but still from one heat) that would be difficult to utilize in any other way.
• We are starting the order of the ceramic tube insulators that make the sextupole’s two water circuit per coil cooling scheme compact enough to fit within the magnet covers.
• DULY Research continued making substantial progress at finishing the Sextupole drawings. A third round of mark-ups (the coils) were sent back to them with minor corrections.
Octupole (OT)
• DULY started and then stopped design of the Octupole while they finish the Sextupole.
Beam Line and Vacuum
• Master Machine started machining the flanges of the Arc Vacuum Chambers and received the material. They are working with a vendor with a 5 axis water jet cutting machine to cut out the parts. We inspected a sample piece of stainless steel that had been water jet cut and determined that there was minimal contamination from the cut method and that in addition, the style of the weld joint would lead to minimal contamination of the vacuum surfaces. We concluded that water jet cutting for these chambers is acceptable.
• For the Arc Chamber job, Master Machine finally received the water jet cut materials. The effect of the delay is about three and a half weeks. They have subcontracted beveling to a Richmond firm.
• Using the GW magnet, the test chambers were tested for their effect on core field and field flatness with the welds shaved. It removed an anomalous spike in one side of the 1/4 inch thick chambers.
• Using the lull in further fabrication of the injection line’s X Chamber (because of missing transition pieces), we tested the effect of the chamber cross section on field. The effect appears to be flat to within the resolution of the hall probe. The chamber welds may not require any machining to flatten the field. We will do one more position to verify the non-effect.
• The Jlab Shop completed welding the X Chamber together and received material to start fabrication of the four chambers for the optical chicane region.
• We examined the inside welds of the X Chamber with a bore scope and found one region of insufficient penetration that required a small re-weld.
• The Jlab Shop is starting fabrication of the four chambers for the optical chicane region. They were then put on hold in order to complete more high priority work.
• The layout of the drawing of the regions around the wiggler was signed off. Design of the girders were started
• The extraction "Y" chamber layout was signed off and the stand positions for the magnets were released to the alignment folks.
• The extraction "Y" chamber work was started in the Jlab Shop.
• The stands for sextupoles and trim quads were fabricated.
• Design of the Octupole stand started
• Girder assembly has started up again with the return of technicians from their CEBAF duties during the latest down period. This time, corrector dipoles that have been characterized and signed off by David Douglas, will be added to the assembly process.
• As the vacuum crew came off CEBAF deployment, girder and vacuum assembly continued with the 5 F girders assembly and installation of the vacuum pipe between cryomodules.
• We issued another revision (L) of the power supply list that included the new requirement for the 4 GC vertical corrector dipoles in the Arcs which will utilize the steel and some of the windings of four QT Trim Quadrupoles. For each of the four systems, a standard trim card power supply is replaced by two low current (2A), power supplies that float with respect to ground and are able to be controlled accurately while run in parallel.
• Low Conductivity Water piping installation continued in the FEL enclosure while additional stand were installed and the alignment crew continued stand placement.
• The first meeting was held to start design of the recirculation dump region.

WBS 11 (Optics):
We implemented a fix for the optical cavity to reduce the binding problems encountered with the Si:bronze/sputtered-silver coated leadscrew. This ameliorated the problem, but did not eliminate it. We did find a component that wasn't made to our drawing tolerances, and fixing that will reduce the problem. Even with the binding problem, the torque of this nut/leadscrew combination was lower (~ 22 in-lbs) than the Vespel/bare stainless steel leadscrew combination (in air).

One of the test/assembly stands was set up with a complete deformable mirror assembly, actuators and LVDT. The assembly went fairly smoothly. We are using this and a portion of the OCMMS test stand to look for low frequency resonances that could prove troublesome when we install the optical cavity hardware in the FEL vault. The good news is that no mechanical resonances attributable to the cavity hardware have been found. This is particularly important for the UV FEL optical cavity, which will utilize an almost identical design. We are using this setup to continue working on the OCMMS feedback circuits. We operated a prototype of the utility (water , power and instrumentation) routing mechanism in excess of 2500 times to see if there were any problems. Some binding and resultant particulate generation was noted, an will be addressed. The contract for the cavity vacuum vessels was awarded. We anticipate receipt near the end (27th) of September.

We have also worked on the OCMMS feedback implementation, and building the support frame for the HEPA filtered enclosure the optical cavity assemblies will be built under when they return from cleaning. The Machine Shop also refinished one of the brackets that connects the leadscrew nut to the rest of the cavity hardware. It was checked on the CMM and is now in spec. Next week we will reassemble and test the hardware with the remanufactured part. We held design meetings to review the mounting brackets for the optical cavity water cooling tubes, and on the optical transport turning cassettes. We also met on the design of the collimator. We received notification that the award for the remaining linear translators and the high vacuum LVDTs were awarded.

We measured the phase noise of our frequency-locked ultrafast laser. When everything is optimized, the value corresponds to a timing jitter of about 400 fs rms. However, there are times when the timing jitter is on the order of 1 ps; too high in many cases to do the measurements we would like to do. After optimizing the alignment of a photodiode that controls the modelocking of the ultrafast laser has improved the power stability and we are now getting consistent values of phase noise due to the PLL. Converted to rms timing jitter, values are no worse than 400 fs and with some judicious tradeoffs (e.g., output power), we can achieve values closer to 250 fs.

We tested the optical cavity mirror motion assembly using a Si/Bronze nut with the sliver-sputtered leadscrew. We found that we still had high drag torque. In part this is due to the silver spalling off the leadscrew. We are preparing to try one of our options; the use of molybdenum disulfide (a JLab-approved UHV lubricant) sputtered onto a bare stainless steel leadscrew. The HEPA filtered enclosure around the optical cavity test and assembly stands was completed.

We found that the use of sprayed-on molybdenum disulfide (a JLab-approved UHV lubricant) onto a bare stainless steel leadscrew resulted in a ~ 2-3x reduction in drag torque when a Si/bronze nut was used. We plan to try sputtering this onto a leadscrew, and are having two silver-plated leadscrews stripped of silver so they can be plated. We are having parts fabricated to do the sputtering in-house, but are investigating whether it can be done by a coating vendor. We also will try tungsten disulfide.

Detailing continues on the optical transport turning cassette and insertable mirror designs. Our outgas rate test stand was rebuilt to test some optical cavity components.

After receiving feedback from the optics design group, we agreed that the optical transport turning cassette design was sound, and it is being detailed. It will be adopted for the UV optical transport as well. As noted earlier, linear translators and LVDTs are already being procured. The collimator design is also going forward. Revisions to the optical cavity vacuum vessels (for the rotary feedthrough and the cooling water jacket) were signed this month.

A few additional parts for the FEL-FROG diagnostic were ordered. Parts for the injector-THz diagnostic are being fabricated. We inspected the mirrors that will be used in the injector light box and they handily met all optical specifications.

Other activities:
We assisted in verifying the performance of a new injector drive laser (for CEBAF), since we are considering purchase of the advanced drive laser from the same vendor. It performed quite nicely. We also spent some time with potential vendors of the laser we'll use with the FEL external cavity (the pulse stacker).

We spent some more time with potential vendors of the laser we'll use with the FEL external cavity (the pulse stacker).

We worked on presentations to be given at the FEL conference and Boulder Damage Conferences. We hosted a visit by Lt. Fulvia Fiorani, NPS, and gave her materials and advice on the construction of a prototype active mirror stabilization system she is building for her Master's thesis. We spent some more time with potential vendors of the laser we'll use with the FEL external cavity (the pulse

Aerospace/Jefferson Lab Subcontract
The preliminary design review for the JLAB laser microengineering station was held on August 22 using a video teleconferencing format that joined The Aerospace Corporation staff with the JLAB staff. Dr. Fred Dylla of JLAB and Dr. John Eric of AFRL co-chaired the review team, which consisted of members of the JLAB FEL staff. Dr. Eric was present in Los Angeles while the JLAB group viewed the presentation via the video linkup. The design review consisted of a formal report and view chart presentation material. The report was sent to JLAB in advance for review. During the three-hour video link-up aspects of the design, for each identified subsystem, were presented and questions were answered. We also presented data on our budget, our spending rate and potential mitigation steps for the expected funding shortfall in FY03. An action item list was generated and we are in the process of working through the list. The overall verbal response of the review was positive and we received a verbal authority to proceed. A written response from JLAB is still pending. The preliminary design review and the supporting material are considered a deliverable and fulfill a contractual obligation.

Other Related Projects

IR Upgrade Commissioning (ONR)
This project was funded in FY02 with $1,000k, with a period of performance from July 1, 2002 to March 30, 2003. Additional funding of approximately $1,000k is anticipated in FY03. In this reporting period, $170k has been expended for a total expenditure of $430k. Commitments total $ 311.6k.
The RF power systems were readied for the first high power tests scheduled in September. The hardware/software interface for re-commissioning the photogun drive laser was prepared for re-commissioning the drive laser in September. The beam physics team continued their iteration on the commissioning plan. The magnetic measurement system for dipoles is fully operational and is being used to characterize the injector dipoles (DV/DU) and optical chicane dipoles (DW) prior to release to the installation crew. The Accelerator Readiness Review ( ARR) was re-instituted in preparation for re-certification of all the FEL systems prior to electron/photon beam generation during the commissioning period

Cryomodule 3
This project was funded in FY02 with $400k. Additional funding of $350k is anticipated in FY03. In this reporting period, $51k has been expended for a total expenditure of $80.8. Commitments total $51.3k.
Cavity cell fabrication and welding continued for assembly of 7-cell structures for the 3rd cryomodule using an improved electron beam welding procedure. The so-called "string assembly" of 8, 7-cell cavity structures for the first CEBAF Upgrade cryomodule was nearly completed and for release to the cryomodule assembly area next month. (This will be the first upgrade cryomodule produced and its assembly and testing precedes the fabrication of the 3rd FEL cryomodule.) Analysis of the first measurements of HOM damping in the 7-cell structure continued. Plans were made to perform additional tests on cold cavities with an appended fundamental power coupler (FPC) to quantify the effects of additional HOM damping provided by the FPC impedance. Also the team is planning to look at the possible advantage of an additional HOM damper placed near the FPC using a test fixture on a copper model.

One Micron Wiggler Project (ONR)
This project was funded in FY02 with $600k. Completion funding of $600k is anticipated in FY03. In this reporting period, $27k has been expended for a total expenditure of $78.7k. Commitments total $17.1k.
A good understanding of the deviations from simple theory when the actual magnetic field of the optical klystron is integrated was developed. This will be presented at the FEL conference in September. The bottom line is that the ideal theory is a good approximation to the exact solution and that, therefore, the spreadsheet model developed previously is a good optimization tool. The spreadsheet indicated that the permanent magnet, uniform wiggler solution is the optimum one. The wiggler wavelength of 6 cm is not commonly used in storage rings. A vendor was contacted on the possibility of using an existing design for the 1 micron wiggler. Some useful solutions were obtained which may be quite cost effective. A design does exist at 5.5 cm wavelength and this was tested and found to be non-optimal but perhaps acceptable. Dave Douglas found an optics solution for this design which works as well as the present solution.

JTO Cryomirror Project
We focused efforts on establishing a firm basis for proceeding with engineering tests of the proposed system. The results discussed under the technical section now gives us confidence in the reliability of the hardware approach proposed.
In this reporting period, $40.8k has been expended for a total expenditure of $60.1. Commitments total $135.9k. The program is on track financially and is expected to deliver all elements of the SOW within projected cost.
We conducted repeated temperature cycling of a sapphire mirror mounted into a nickel-plated molybdenum holder.  The mounting was done using our improved "advanced braze" technique.  The mirror was cooled at far faster rates than planned,  up to > 40 K/min, in an effort to test the durability of the braze.  After ~ 10 cyclings, the mirror is still solidly mounted.

JTO Drive Laser Project
After verifying the claimed performance versus vendor claims we are in the process of proceeding on the major laser hardware purchase.
In this reporting period, $11.7k has been expended for a total expenditure of $19.5k. Commitments total $4.5k. The program is on track financially and is expected to deliver all elements of the SOW within projected cost.
We measured the timing jitter of a custom Time-Bandwidth laser (our lead candidate for the drive laser front end) that was purchased by the Accelerator Division's Injector Group.  The laser's timing jitter was ~ 0.5 ps, within our spec.  Discussions with company representatives give us confidence that we may do even better with the laser they would build for us. We requested an updated quote from them, and are ready to go forward with a purchase request on receipt of that updated quote.